# Evidence of Energy Injection in the Short and Distant GRB 250221A

### v4

@@ Abstract
We present the photometric and spectroscopic analysis of the short-duration GRB 250221A ($T_{90}=1.80\pm 0.32$ s), using a data set from the optical facilities COLIBRÍ, the Harlingten 50 cm Telescope, and the Very Large Telescope. We complement these observations with data from the Neil Gehrels Swift Observatory and the Einstein Probe, as well as radio observations from the Very Large Array. GRB 250221A is among the few short GRBs with direct afterglow spectroscopy, which gives a secure redshift determination of $z=0.768$ and allows the unambiguous identification of the host as a galaxy with a star-formation rate of $\sim 3\,M_{\odot}\,{\rm yr}^{-1}$. The X-ray and optical light curves up to $T_{0}+10$ ks (where $T_{0}$ refers to the GRB trigger time) are well described by forward-shock synchrotron emission in the slow-cooling regime within the standard fireball framework. However, at $T_{0}+0.6$ days, both the X-ray and optical bands exhibit an excess over the same interval, which we interpret as evidence of energy injection into a jet with a half-opening angle of $\theta_{j}=11.5^{\circ}$ through a refreshed shock powered by late central engine activity or a radially stratified ejecta. The burst properties (duration, spectral hardness, peak energy, and location in the Amati plane) all favour a compact binary merger origin. However, our modelling of the afterglow suggests a dense circumburst medium ($n\sim 80$ cm-3), which is more typical of a Collapsar environment. This tension over the classification of this burst (short-hard vs. long-soft) as inferred from the prompt and afterglow emissions makes GRB 250221A an unusual event and underscores the limitations of duration-based classifications and the importance of multi-wavelength, time-resolved follow-up observations.
@@
pubyear: 2025pagerange: Evidence of Energy Injection in the Short and Distant GRB 250221AB
@@ 1 Introduction
Gamma-ray bursts (GRBs) are the most luminous electromagnetic phenomena observed in the universe (Atteia et al., 2017) and are powered by ultra-relativistic jets, with an overall duration dichotomy that is usually associated to two different types of progenitors (Kouveliotou et al., 1993). Short GRBs (SGRBs) have durations $T_{90}$ [^1] The duration over which the central 90% of the photons from the GRB are detected. $\lesssim 2$ seconds and a relatively hard spectrum, characterized through the hardness ratio (HR [^2] The energy or fluence contained within one band compared to another, where SGRBs usually emit more high-energy radiation than LGRBs.), and are believed to result from the merger of compact binary systems with at least one neutron star (NS) involved (see, e.g., Paczynski, 1986; Paczyński, 1991; Lee & Ramirez-Ruiz, 2007; Berger, 2014; Abbott et al., 2017). On the other hand, long GRBs (LGRBs) are thought to be the result of the core-collapse of a Wolf-Rayet star (with a stripped H and He envelope) whose mass exceeds about $10M_{\odot}$ and are associated with type Ic supernovae (SNe) (see, e.g., Woosley, 1993; MacFadyen & Woosley, 1999a; Hjorth & Bloom, 2012).
@@
However, this duration-based classification has important limitations, which can be seen in several cases, including: hybrid events such as the long GRB 211211A ($T_{90}=51.37\pm 0.80$s) (Troja et al., 2022; Yang et al., 2022; Rastinejad et al., 2022) and GRB 230307A ($T_{90}\sim 35$s) (Yang et al., 2024; Levan et al., 2024), but with an identified kilonovae (KNe) signature; the short-duration GRB 200826A ($T_{90}=1.4\pm 0.13$s) linked to the collapse of a massive star (Ahumada et al., 2021); and unusual events such as GRB 210704A possibly produced by the coalescence between a neutron star and a white dwarf (Becerra et al., 2023). This motivates the use of additional diagnostics, in particular the properties of the afterglow and host environment, to distinguish between progenitors.
@@
Supernovae and kilonovae are the smoking gun signatures of the progenitors. However, they are not detectable at high redshift. In such cases, the afterglow and burst environment become the primary diagnostics. In this context, the multi-frequency afterglow is a powerful probe. While the fireball model (Goodman, 1986; Paczynski, 1986; Shemi & Piran, 1990; Rees & Meszaros, 1992; Meszaros & Rees, 1993), captures the overall prompt and afterglow evolution (see Kumar & Zhang, 2015, for a review), many optical light curves show deviations such as flares, plateaus, and rebrightening episodes (Sari & Mészáros, 2000; Lazzati et al., 2002; Petropoulou et al., 2020; Dainotti et al., 2024; Gendre, 2025). The physical origin of these features remains debated, with possible explanations including late-time central engine activity (Li et al., 2012), energy injection (Laskar et al., 2015), refreshed shocks (Vlasis et al., 2011; Lamb et al., 2019), or interaction with a structured circumburst medium (CBM). Such signatures are particularly relevant when the progenitor classification is uncertain, since they may reveal information about both the central engine and the environment.
Complementary to the photometric information, spectroscopy is essential for a complete understanding of every event. It provides the only direct means of determining redshift, offers clues about the host galaxy and the environment of the burst, and helps constrain the progenitor system.
@@
Optical spectroscopy of SGRB host galaxies is particularly challenging because these systems are typically faint and associated with older stellar populations (O’Connor et al., 2022; Fong et al., 2022; Nugent et al., 2022). Nonetheless, when successful, such observations yield valuable information on the global environment in which these transients occur.
@@
In this paper, we present an analysis of both the early- and late-time observations of the multi-frequency afterglow of GRB 250221A. This is one of the few SGRB afterglows (only the fourth case in the sample of over 150 SGRBs discovered by Swift) with clear absorption lines in its spectrum, allowing us to securely measure its redshift $z\!\approx\!0.768$ and break the degeneracy with a nearby, unrelated galaxy at $z\!\approx\!0.343$. Its late-time afterglow shows an unusual rebrightening starting at $\sim$0.6 d post-burst — later and more pronounced than in most previously reported events (see, e.g., Troja et al., 2007; Becerra et al., 2019a, b; Pereyra et al., 2022; Nardini et al., 2011; Kobayashi & Zhang, 2003; Li et al., 2012; Kann et al., 2010) — making it an outstanding event for investigating how such features connect to progenitor type and environment. By combining early- and late-time multi-wavelength photometry with spectroscopy of the host candidates, this burst allows us to explore the mechanisms behind rebrightening and assess their implications for GRB classification. We examine its high-energy characteristics during the prompt phase, including duration, isotropic energy, and spectral peak energy, along with multi-wavelength observations of the afterglow in the X-ray, optical, and radio bands extending up to 10 days post-burst.
@@ 2.2.2 Harlingten 50-cm Telescope
Figure 1: Top: X-ray (0.3–10.0 keV, at 1 keV, black circles), optical (yellow diamonds for i, red hexagons for r, green stars for g) and radio (at 9 GHz from ATCA (Gulati et al., 2025) and 10 GHz from VLA, brown pentagons) observations of GRB 250221A. The dashed lines show the power-law fits to the afterglow for X-rays (fitted before $T\lesssim 2\times 10^{4}$ s and after $T>10^{5}$ s) and optical (fitted for $T<10^{4}$ s for the i band and at $T>10^{5}$ s for the late r band). Bottom: Optical-to-X-rays spectral indices for three different epochs, the first from $T_{0}+100$ s to $T_{0}+10^{4}$ s, with a median arrival time of the X-rays photons at $T_{0}+1800$ s (see Figure 7) (purple), second at $T_{0}+9\times 10^{4}$ s (blue), and third at $T_{0}+1.8\times 10^{5}$ s (green). We also illustrate the optical light curves of SN1998bw (Galama et al., 1998; Clocchiatti et al., 2011) and SN2011kl (Greiner et al., 2015) (placed at $z=0.768$) to highlight that the excess is similar in brightness (in the optical) but differs in timescale.
@@ 2.2.3 Very Large Telescope
Figure 2: Field of GRB 250221A. We show the R filter image obtained with the VLT/FORS2 after $T_{0}+0.9$ days (panel a) and $T_{0}+11.9$ days (panel b). We also show the contribution of the afterglow after image subtraction (panel c). We highlight the presence of two galaxies at $photo-z=0.343\pm 0.077$ (G1, green circle; according to the Legacy Survey DR10 Dey et al., 2019) and at $z=0.768$ (G2, magenta circle).
@@ 2.2.4 Other Optical and Infrared Observations
Optical observations of GRB 250221A were also performed by the Rapid Eye Mount (REM) 60 cm robotic telescope (Melandri et al., 2025), 0.36 m telescope in the Al-Khatim Observatory (Odeh et al., 2025), 1.6 m Multi-channel Photometric Survey Telescope (Mephisto) (Guo et al., 2025), the Nordic Optical Telescope (NOT) Cotter et al. (2025), GRANDMA/TAROT (Muenter et al., 2025), the AZT-20 telescope at the Assy-Turgen Observatory (Pankov et al., 2025a), the AZT-33IK telescope at the Mondy Observatory (Pankov et al., 2025b), the 1-meter Sinistro telescope at the Las Cumbres Observatory Global Telescope (LCOGT) (Ghosh et al., 2025), the 80 cm Tsinghua-Nanshan Optical Telescope (TNOT) (Iskandar et al., 2025). GRB 250221A was also detected in the near IR, in the $H$ filter, by REM (Melandri et al., 2025).
Table 1: Optical and radio photometry of GRB 250221A from this work. Optical values are corrected by Galactic extinction.
@@
Telescope Instrument Mid Time [hrs] Energy Range Exposure Time [ks] Flux [erg cm-2]
EP FXT 49.68 0.3–10 5 $(2.28_{-0.24}^{+0.28})\times 10^{-13}$
EP FXT 136.32 0.3–10 4 $(2.36_{-1.02}^{+1.35})\times 10^{-14}$
@@
COLIBRÍ DDRAGO 0.03 i 16.80 $\pm$ 0.04 60 $634.08\pm 21.35$
COLIBRÍ DDRAGO 0.05 i 17.41 $\pm$ 0.05 60 $360.36\pm 18.00$
COLIBRÍ DDRAGO 0.08 i 17.72 $\pm$ 0.07 60 $270.02\pm 16.68$
COLIBRÍ DDRAGO 0.12 i 18.04 $\pm$ 0.11 60 $200.36\pm 20.93$
COLIBRÍ DDRAGO 0.14 i 18.30 $\pm$ 0.08 60 $158.09\pm 11.95$
COLIBRÍ DDRAGO 0.16 i 18.52 $\pm$ 0.06 60 $129.36\pm 7.12$
COLIBRÍ DDRAGO 0.18 i 18.65 $\pm$ 0.08 60 $114.73\pm 8.11$
COLIBRÍ DDRAGO 0.21 i 18.73 $\pm$ 0.08 60 $106.87\pm 7.56$
COLIBRÍ DDRAGO 0.23 i 18.73 $\pm$ 0.08 60 $106.59\pm 7.49$
COLIBRÍ DDRAGO 0.31 i 19.02 $\pm$ 0.09 300 $17.89\pm 6.81$
COLIBRÍ DDRAGO 0.42 i 19.32 $\pm$ 0.19 300 $81.54\pm 6.78$
COLIBRÍ DDRAGO 0.53 i 19.70 $\pm$ 0.28 300 $62.17\pm 11.08$
COLIBRÍ DDRAGO 0.64 i 19.92 $\pm$ 0.10 300 $43.60\pm 11.29$
COLIBRÍ DDRAGO 0.83 i 20.12 $\pm$ 0.11 300 $35.63\pm 3.40$
COLIBRÍ DDRAGO 0.94 i 20.22 $\pm$ 0.25 300 $29.57\pm 2.97$
COLIBRÍ DDRAGO 1.29 g 20.69 $\pm$ 0.25 300 $27.15\pm 6.19$
COLIBRÍ DDRAGO 1.32 r 20.64 $\pm$ 0.41 300 $16.00\pm 4.00$
COLIBRÍ DDRAGO 1.34 i 20.90 $\pm$ 0.31 300 $14.50\pm 4.20$
50cm Harlingten 7.80 r 21.67 $\pm$ 0.52 960 $6.92\pm 3.31$
VLT FORS2 21.12 i 21.91 $\pm$ 0.10 720 $4.37\pm 0.49$
VLT FORS2 21.12 r 22.17 $\pm$ 0.06 720 $5.70\pm 0.57$
COLIBRÍ DDRAGO 23.85 r 21.94 $\pm$ 0.27 5640 $5.46\pm 1.35$
COLIBRÍ DDRAGO 47.85 r 21.63 $\pm$ 0.24 5160 $7.15\pm 1.61$
COLIBRÍ DDRAGO 72.21 r 22.28 $\pm$ 0.25 4020 $3.96\pm 0.93$
VLT FORS2 117.12 r 23.36 $\pm$ 0.07 1440 $1.46\pm 0.41$
VLT FORS2 117.60 i 23.00 $\pm$ 0.15 720 $2.08\pm 0.31$
@@
VLA 46.80 X (10 GHz) 900 $63.00\pm 9.00$
VLA 238.80 X (10 GHz) 900 $200.00\pm 12.00$

### v5

@@ Abstract
We present the photometric and spectroscopic analysis of the short-duration GRB 250221A ($T_{90}=1.80\pm 0.32$ s), using a data set from the optical facilities COLIBRÍ, the Harlingten 50 cm Telescope, and the Very Large Telescope. We complement these observations with data from the Neil Gehrels Swift Observatory and the Einstein Probe, as well as radio observations from the Very Large Array. GRB 250221A is among the few short GRBs with direct afterglow spectroscopy, which gives a secure redshift determination of $z=0.768$ and allows the unambiguous identification of the host as a galaxy with a star-formation rate of $\sim 3\,M_{\odot}\,{\rm yr}^{-1}$. The X-ray and optical light curves up to $T_{0}+3\times 10^{4}$ s (where $T_{0}$ refers to the GRB trigger time) are well described by forward-shock synchrotron emission in the slow-cooling regime within the standard fireball framework. However, at $T_{0}\sim 5\times 10^{4}$ s, both the X-ray and optical bands exhibit an excess over the same interval, which we interpret as evidence of energy injection into a jet with a half-opening angle of $\theta_{j}=11.5^{\circ}$ through a refreshed shock powered by late central engine activity or a radially stratified ejecta. The burst properties (duration, spectral hardness, peak energy, and location in the Amati plane) all favour a compact binary merger origin. However, our modelling of the afterglow suggests a dense circumburst medium ($n\sim 80$ cm-3), which is more typical of a collapsar environment.
@@
pubyear: 2026pagerange: Evidence of Energy Injection in the Short and Distant GRB 250221A in a High Density EnvironmentB
@@ 1 Introduction
Gamma-ray bursts (GRBs) are the most luminous electromagnetic phenomena observed in the universe (Atteia et al., 2017) and are powered by ultra-relativistic jets, with an overall duration dichotomy (Kouveliotou et al., 1993) that is usually associated to two different types of progenitors. Short GRBs (SGRBs) have durations $T_{90}$ [^1] The duration over which the central 90% of the photons from the GRB are detected. $\lesssim 2$ seconds and a relatively hard spectrum, characterized through the hardness ratio (HR [^2] The energy or fluence contained within one band compared to another, where SGRBs usually emit more high-energy radiation than LGRBs.), and are believed to result from the merger of compact binary systems with at least one neutron star (NS) involved (see, e.g., Paczynski, 1986; Paczyński, 1991; Lee and Ramirez-Ruiz, 2007; Berger, 2014; Abbott et al., 2017). On the other hand, long GRBs (LGRBs) are thought to be the result of the core-collapse of a Wolf-Rayet star (with a stripped H and He envelope) whose mass exceeds about $10M_{\odot}$ and are associated with type Ic supernovae (SNe) (see, e.g., Woosley, 1993; MacFadyen and Woosley, 1999b; Hjorth and Bloom, 2012).
@@
However, this duration-based classification has important limitations, which can be seen in several cases, including: hybrid events such as the long GRB 211211A ($T_{90}=51.37\pm 0.80$s) (Troja et al., 2022; Yang et al., 2022; Rastinejad et al., 2022) and GRB 230307A ($T_{90}\sim 35$s) (Yang et al., 2024; Levan et al., 2024), but with an identified kilonovae (KNe) signature; the short-duration GRB 200826A ($T_{90}=1.4\pm 0.13$s) linked to the collapse of a massive star (Ahumada et al., 2021); and unusual events such as GRB 210704A likely produced by the coalescence between a neutron star and a white dwarf (Becerra et al., 2023), combining properties exhibited by collapsar and merger-driven bursts. This motivates the use of additional diagnostics, in particular the properties of the afterglow and host environment, to distinguish between progenitors.
Beyond duration and hardness, several studies have therefore explored more physically motivated classification schemes. In particular, the type-I/type-II framework (e.g., Zhang, 2007) incorporates host-galaxy characteristics, energetics, and afterglow behaviour to differentiate merger-driven from collapsar-driven events. This broader approach has also been applied by Tsvetkova et al. (2017) when classifying short GRBs, underscoring the need for a multidimensional view of GRB origins.
@@
Supernovae and kilonovae provide the most direct evidence of the underlying progenitors. However, they are not detectable at high redshift. In such cases, the afterglow and burst environment become the primary diagnostics. In this context, the multi-frequency afterglow is a powerful probe. While the fireball model (Goodman, 1986; Paczynski, 1986; Shemi and Piran, 1990; Rees and Meszaros, 1992; Meszaros and Rees, 1993), captures the overall prompt and afterglow evolution (see Kumar and Zhang, 2015; Agüí Fernández et al., 2023, for a review), many optical light curves show deviations such as flares, plateaus, and rebrightening episodes (Sari and Mészáros, 2000; Lazzati et al., 2002; Petropoulou et al., 2020; Dainotti et al., 2024; Gendre, 2025). The physical origin of these features remains debated, with possible explanations including late-time central engine activity (Li et al., 2012), energy injection (Laskar et al., 2015), refreshed shocks (Vlasis et al., 2011; Lamb et al., 2019), or interaction with a structured circumburst medium (CBM). Such signatures are particularly relevant when the progenitor classification is uncertain, since they may reveal information about both the central engine and the environment.
@@
Complementary to the photometric information, spectroscopy is essential for a complete understanding of every event. It provides the only direct means of determining redshift, offers clues about the host galaxy and the environment of the burst, and helps constrain the progenitor system. Optical spectroscopy of SGRB host galaxies is particularly challenging because these systems are typically faint and associated with older stellar populations (O’Connor et al., 2022; Fong et al., 2022; Nugent et al., 2022). Nonetheless, when successful, such observations yield valuable information on the global environment in which these transients occur.
@@
In this paper, we present an analysis of both the early- and late-time observations of the multi-frequency afterglow of GRB 250221A. This is one of the few SGRB afterglows (only the fourth case in the sample of over 150 SGRBs discovered by Swift) with clear absorption lines in its spectrum, allowing us to securely measure its redshift $z\!\approx\!0.768$ and break the degeneracy with a nearby, unrelated galaxy at $z\!\approx\!0.343$. Its late-time afterglow shows an unusual rebrightening starting at $\sim$0.6 d post-burst — later and more pronounced than in most previously reported events (see, e.g., Troja et al., 2007; Becerra et al., 2019b, a; Pereyra et al., 2022; Nardini et al., 2011; Kobayashi and Zhang, 2003; Li et al., 2012; Kann et al., 2010) — making it an outstanding event for investigating how such features connect to progenitor type and environment. By combining early- and late-time multi-wavelength photometry with spectroscopy of the host candidates, this burst allows us to explore the mechanisms behind rebrightening and assess their implications for GRB classification. We examine its high-energy characteristics during the prompt phase, including duration, isotropic energy, and spectral peak energy, along with multi-wavelength observations of the afterglow in the X-ray, optical, and radio bands extending up to 10 days post-burst.
@@ 2.2.2 Harlingten 50-cm Telescope
Figure 1: Top: X-ray (0.3–10.0 keV, at 1 keV, black circles), optical (yellow diamonds for i, red hexagons for r, green stars for g) and radio (at 9 GHz from ATCA (Gulati et al., 2025) and 10 GHz from VLA, brown pentagons) observations of GRB 250221A. The dashed lines show the power-law fits to the afterglow for X-rays (fitted before $T\lesssim 2\times 10^{4}$ s and after $T>10^{5}$ s) and optical (fitted for $T<10^{4}$ s for the i band and at $T>10^{5}$ s for the late r band). Bottom: Optical-to-X-rays spectral indices for three different epochs, the first from $T_{0}+100$ s to $T_{0}+10^{4}$ s, with a median arrival time of the X-rays photons at $T_{0}+1800$ s (see Figure 7 and Section 4.2) (purple), second at $T_{0}+9\times 10^{4}$ s (blue), and third at $T_{0}+1.8\times 10^{5}$ s (green). We also illustrate the optical light curves of SN1998bw (Galama et al., 1998; Clocchiatti et al., 2011) and SN2011kl (Greiner et al., 2015) (placed at $z=0.768$) to highlight that the excess is similar in brightness (in the optical) but differs in timescale.
@@ 2.2.3 Very Large Telescope
Figure 2: Field of GRB 250221A. We show the R filter image obtained with the VLT/FORS2 after $T_{0}+0.9$ days (panel a) together with the position of the X-Shooter slit used to acquire the spectrum (orange box)
and $T_{0}+11.9$ days (panel b). We also show the contribution of the afterglow after image subtraction (panel c). We highlight the presence of two galaxies at $photo-z=0.343\pm 0.077$ (G1, green circle; according to the Legacy Survey DR10 Dey et al., 2019) and at $z=0.768$ (G2, magenta circle).
@@ 2.2.4 Other Optical and Infrared Observations
Optical observations of GRB 250221A were also performed by the Rapid Eye Mount (REM) 60 cm robotic telescope (Melandri et al., 2025), 0.36 m telescope in the Al-Khatim Observatory (Odeh et al., 2025), 1.6 m Multi-channel Photometric Survey Telescope (Mephisto) (Guo et al., 2025), the Nordic Optical Telescope (NOT) Cotter et al. (2025), GRANDMA/TAROT (Muenter et al., 2025), the AZT-20 telescope at the Assy-Turgen Observatory (Pankov et al., 2025a), the AZT-33IK telescope at the Mondy Observatory (Pankov et al., 2025b), the 1-meter Sinistro telescope at the Las Cumbres Observatory Global Telescope (LCOGT) (Ghosh et al., 2025), the 80 cm Tsinghua-Nanshan Optical Telescope (TNOT) (Iskandar et al., 2025). GRB 250221A was additionally detected in the near IR by REM (Melandri et al., 2025).
These observations are reported here for completeness; however, they are not used in the analysis presented in this work, as they do not provide additional constraints on the properties discussed below.
Table 1: Optical and radio photometry of GRB 250221A used in this work. Optical values are corrected by Galactic extinction.
@@
[…]
Swift XRT 149.82 0.3–10 $9.24\pm 2.02$ This work
Swift XRT 173.06 0.3–10 $8.01\pm 1.79$ This work
Swift XRT 204.19 0.3–10 $9.72\pm 2.08$ This work
Swift XRT 242.75 0.3–10 $4.99\pm 1.12$ This work
Swift XRT 309.46 0.3–10 $3.61\pm 0.81$ This work
Swift XRT 360.39 0.3–10 $4.98\pm 1.09$ This work
Swift XRT 422.20 0.3–10 $3.80\pm 0.84$ This work
Swift XRT 471.64 0.3–10 $5.58\pm 1.26$ This work
Swift XRT 554.48 0.3–10 $3.26\pm 0.53$ This work
Swift XRT 5607.76 0.3–10 $0.36\pm 0.04$ This work
Swift XRT 23784.89 0.3–10 $0.05\pm 0.01$ This work
Swift XRT 78038.44 0.3–10 $0.03\pm 0.01$ This work
EP FXT 178848.0 0.3–10 5000 $(2.82_{-0.27}^{+0.29})\times 10^{-2}$ This work
EP FXT 490752.0 0.3–10 4000 $(2.91_{-0.13}^{+0.16})\times 10^{-2}$ This work
@@
COLIBRÍ DDRAGO 108.0 i 16.80 $\pm$ 0.04 60 $634.08\pm 21.35$ This work
COLIBRÍ DDRAGO 180.0 i 17.41 $\pm$ 0.05 60 $360.36\pm 18.00$ This work
COLIBRÍ DDRAGO 288.0 i 17.72 $\pm$ 0.07 60 $270.02\pm 16.68$ This work
COLIBRÍ DDRAGO 432.0 i 18.04 $\pm$ 0.11 60 $200.36\pm 20.93$ This work
COLIBRÍ DDRAGO 504.0 i 18.30 $\pm$ 0.08 60 $158.09\pm 11.95$ This work
COLIBRÍ DDRAGO 576.0 i 18.52 $\pm$ 0.06 60 $129.36\pm 7.12$ This work
COLIBRÍ DDRAGO 648.0 i 18.65 $\pm$ 0.08 60 $114.73\pm 8.11$ This work
COLIBRÍ DDRAGO 756.0 i 18.73 $\pm$ 0.08 60 $106.87\pm 7.56$ This work
COLIBRÍ DDRAGO 828.0 i 18.73 $\pm$ 0.08 60 $106.59\pm 7.49$ This work
COLIBRÍ DDRAGO 1116.0 i 19.02 $\pm$ 0.09 300 $17.89\pm 6.81$ This work
COLIBRÍ DDRAGO 1512.0 i 19.32 $\pm$ 0.19 300 $81.54\pm 6.78$ This work
COLIBRÍ DDRAGO 1908.0 i 19.70 $\pm$ 0.28 300 $62.17\pm 11.08$ This work
COLIBRÍ DDRAGO 2304.0 i 19.92 $\pm$ 0.10 300 $43.60\pm 11.29$ This work
COLIBRÍ DDRAGO 2988.0 i 20.12 $\pm$ 0.11 300 $35.63\pm 3.40$ This work
COLIBRÍ DDRAGO 3384.0 i 20.22 $\pm$ 0.25 300 $29.57\pm 2.97$ This work
COLIBRÍ DDRAGO 4644.0 g 20.69 $\pm$ 0.25 300 $27.15\pm 6.19$ This work
COLIBRÍ DDRAGO 4752.0 r 20.64 $\pm$ 0.41 300 $16.00\pm 4.00$ This work
COLIBRÍ DDRAGO 4824.0 i 20.90 $\pm$ 0.31 300 $14.50\pm 4.20$ This work
50cm Harlingten 28080.0 r 21.67 $\pm$ 0.52 960 $6.92\pm 3.31$ This work
VLT FORS2 76032.0 i 21.91 $\pm$ 0.10 720 $4.37\pm 0.49$ This work
VLT FORS2 76032.0 r 22.17 $\pm$ 0.06 720 $5.70\pm 0.57$ This work
COLIBRÍ DDRAGO 85860.0 r 21.94 $\pm$ 0.27 5640 $5.46\pm 1.35$ This work
COLIBRÍ DDRAGO 172260.0 r 21.63 $\pm$ 0.24 5160 $7.15\pm 1.61$ This work
COLIBRÍ DDRAGO 259955.1 r 22.28 $\pm$ 0.25 4020 $3.96\pm 0.93$ This work
VLT FORS2 421632.0 r 23.36 $\pm$ 0.07 1440 $1.46\pm 0.41$ This work
VLT FORS2 423360.0 i 23.00 $\pm$ 0.15 720 $2.08\pm 0.31$ This work
@@
ATCA 360.0 9 GHz 36000 <24<24 Gulati et al. (2025)
VLA 168480.0 X (10 GHz) 900 $63.00\pm 9.00$ This work
ATCA 190080.0 9 GHz 14400 $233.00\pm 15.00$ Gulati et al. (2025)
VLA 859680.0 X (10 GHz) 900 $200.00\pm 12.00$ This work